Hi
an inground swimming pool is cylindrical. It has a height of 1.2m radius 1.8m. The floor and walls are to be cement rendered. Calculate the area to be rendered.
My answer is 21.49sq m textbook says its 23.75sq m
thanks

Answer by rothauserc(4718)

surface area(SA) = 2 × pi × r2 + 2 × pi × r × h
:
we do not include the area of the circle on the surface of the pool since it is water :-)
Rendered area(RA) = (pi × r^2) + (2 × pi × r × h)
:
*******************************************************
RA = (22÷7)×((1.8)^2)+2×(22÷7)×(1.8)×(1.2) = 23.76 sq m
*******************************************************
: